Cobalt crystallizes in the hexagonal system (hcp) with the following unit cell parameters a = 2.51Å et c=4.07 Å. Give the radius of cobalt atoms (introduce the result in Å and with three significant figures)
Answer #1

A unit cell characterized by six parameters. These parameters are three edges (a,b and c) and angel between them (alfa, beta and gama).

Radius r=a/2

Radius r =2.51/2 =1.255 A